# QGait: Toward Accurate Quantization for Gait Recognition with Binarized Input

### v1

@@ Abstract
Existing deep learning methods have made significant progress in gait recognition. Typically, appearance-based models binarize inputs into silhouette sequences. However, mainstream quantization methods prioritize minimizing task loss over quantization error, which is detrimental to gait recognition with binarized inputs. Minor variations in silhouette sequences can be diminished in the network’s intermediate layers due to the accumulation of quantization errors. To address this, we propose a differentiable soft quantizer, which better simulates the gradient of the round function during backpropagation. This enables the network to learn from subtle input perturbations. However, our theoretical analysis and empirical studies reveal that directly applying the soft quantizer can hinder network convergence. We further refine the training strategy to ensure convergence while simulating quantization errors. Additionally, we visualize the distribution of outputs from different samples in the feature space and observe significant changes compared to the full precision network, which harms performance. Based on this, we propose an Inter-class Distance-guided Distillation (IDD) strategy to preserve the relative distance between the embeddings of samples with different labels. Extensive experiments validate the effectiveness of our approach, demonstrating state-of-the-art accuracy across various settings and datasets. The code will be made publicly available.
@@ 1 Introduction
In recent years, gait recognition has emerged as a prominent field due to its potential in long-range pedestrian retrieval based on walking patterns. Unlike other biometric technologies, gait, as a form of locomotion, is difficult to disguise [18; 8]. Additionally, gait recognition is robust against common covariates such as attire, carrying items, and standing conditions [13; 7]. Therefore, gait recognition methods hold significant promise for capturing biometric features remotely in uncontrolled environments, thus offering broad application prospects.
@@
With the development of deep learning, gait recognition methods based on deep features [38; 27; 12; 15; 22] have achieved notable results on various datasets [40; 33; 43; 31; 46; 11; 10]. Currently, state-of-the-art gait recognition technologies can achieve recognition distances exceeding hundreds of meters even with 4K high-definition cameras [31]. However, existing methods have scarcely considered the increasing computational resources required by the advanced devices. In some scenarios, it may be necessary to test tens of thousands of subjects within a short timeframe. While methods based on CNNs [11; 43] or Transformers [10] have demonstrated excellent performance, edge devices in deployment environments cannot meet the demands of these resource-intensive algorithms. This necessitates reducing the memory and computational burden of gait recognition methods while preserving performance, enabling deployment on resource-constrained devices. Current efforts to compress and accelerate neural networks primarily focus on quantization [45; 6; 9; 19] and pruning [24; 44; 14]. Compared to pruning, quantization is more deployment-friendly and offers significant acceleration, particularly Quantization Aware Training [16].
@@
Existing quantization methods [42; 1; 9] have achieved decent results on some tasks, but applying them to gait recognition tasks has failed to deliver comparable performance. These methods use a Straight-Through Estimator (STE) [2] to approximate the gradient of non-differentiable operations during the backward pass. While STE can smoothly transfer gradients, it is not sensitive to decimal-level changes. This may have no impact on other tasks but can be fatal for changes in gait posture. For example, in RGB input, the variation from 0 to 1 in a small area is not significant. However, for binarized gait silhouettes, this may indicate a change in posture. In this paper, we propose using parameterized soft quantization operators to prevent the network from accumulating quantization errors. Initially, we attempted to directly utilize a soft quantizer in the training process as a replacement for STE. However, through theoretical analysis and empirical validation, we have demonstrated that although this approach can simulate the quantization error introduced by non-differentiable operations, it significantly impacts the network’s convergence. To harness the strengths of both STE and the soft quantizer while mitigating their respective limitations, we propose a two-stage training strategy. In the first stage, we employ STE to facilitate the network’s convergence towards the vicinity of the optimal solution. Subsequently, we utilize the soft quantizer in the fine-tuning stage, enabling the simulation of decimal-level errors, which improves the network performance.
@@
After quantization, the 8-bit models exhibit outstanding performance, surpassing even the full-precision model. However, when quantized to lower bit widths, the model’s accuracy experiences significant degradation. Through analysis of the sample distribution in feature space, we observed significant changes in the relative positions of samples between the 4-bit quantized model and the full-precision model. This change could be a significant contributing factor to the observed decrease in accuracy. This insight prompts us to consider whether knowledge distillation could be employed to transfer the knowledge from the full-precision model to the low-bit model. However, through an examination of the properties of traditional distillation [9; 17; 39], we found that such methods result in the low-bit model learning the numerical discrepancies between the outputs of the full-precision model and itself, which are discrepancies inherently introduced by quantization. Experimental results similarly demonstrate that this distillation approach does not yield benefits. Therefore, we propose an Inter-class Distance-guided Distillation method, allowing the network to focus more on the differences in sample distribution between classes rather than the numerical differences within classes. This distillation approach leads to a significant improvement in accuracy.
@@
We introduce, for the first time, a quantization-based compressed gait recognition network, which performs lossless compression and acceleration on existing methods as Fig. 1 shows.
We propose a two-stage quantization training strategy together with the IDD method, which simultaneously optimizes quantization error and task loss.
@@ 2.1 Gait Recognition
Gait recognition methods can generally be classified into two categories: model-based methods [20; 21] and appearance-based methods [37; 32]. Model-based approaches tend to utilize estimated human body models as input, such as 2D/3D poses and SMPL [26]. For instance, GaitGraph [34] employs graph convolutional networks for learning gait representations based on 2D skeletons, SMPLGait [43] leverages the 3D geometric information of SMPL models to enhance the learning of gait appearance features, and GaitTR [41] combines transformer and convolutional layers to represent spatial and temporal information respectively. However, although model-based methods theoretically exhibit robustness to factors like carrying items and clothing, they often perform poorly due to the error accumulation in the pre-processing stage while entailing high computational costs, thereby potentially lacking practicality in certain real-world scenarios. On the other hand, appearance-based methods directly learn shape features from input videos, offering simplicity and ease of use while preserving privacy. With the rise of deep learning, most current appearance-based works focus on spatial feature extraction and gait temporal modeling. Specifically, GaitSet [4] treats gait sequences as a set for the first time, employing the maximum function to compress sequences of frame-level spatial features. GaitPart [12] meticulously explores minor differences in various parts of the silhouette to model periodic gait features. GaitGL [22] addresses the limitations of spatially global gait representations in neglecting details, and locally based descriptors in failing to capture relationships between adjacent parts, by developing global and local convolutional layers. Building upon these works, GaitBase [11] significantly simplifies gait modeling and achieves excellent results with its simple yet effective design. Considering that existing methods often lack a practical perspective, it becomes necessary to design a lightweight network to aid gait recognition models in real-world applications. To the best of our knowledge, the proposed QGait is the first quantization-based compressed gait recognition network.
@@ 2.2 Network Quantization
Quantization [5; 3; 28] is a widely adopted technique in computer vision for compressing and accelerating neural networks. Depending on the mapping strategy, quantization can be categorized into uniform and non-uniform [25] quantization. Non-uniform quantization necessitates specific hardware support, thus current research predominantly focuses on uniform quantization exploration. Uniform quantization methods encompass Post-Training Quantization (PTQ) [30] and Quantization Aware Training (QAT). PTQ has gained popularity for its ability to quantize models without requiring retraining. However, it suffers from accuracy degradation due to its heavy reliance on input data for setting quantization parameters, leading to poor alignment between parameters and weights. QAT addresses this issue by training with inserted fake quantization nodes [16], allowing network weights to adapt to quantization errors and enhancing the accuracy of quantized models. Early works like Dorefa [45] achieve network acceleration by quantizing activation values, weights, and gradients simultaneously. Further advancements such as PACT [6], utilize learnable parameters to clip activation values to gain more reasonable quantization intervals. LSQ [9], one of the most widely used quantization methods, introduces a novel approach to estimate and scale the task loss gradient on the quantizer step size for each weight and activation layer, enabling the quantization step size to be learned with network optimization and achieving commendable results across various tasks. Given the satisfactory accuracy and notable acceleration achieved by quantized networks, quantization-based model compression has been extensively studied in various domains [35; 39; 23], demonstrating the efficacy of quantization schemes.
@@ 3.1 Preliminaries
where denotes the embeddings, is the separate fully connected layer, is the temporal pooling and horizontal pooling layer, is the backbone network, and denotes the connection among network parts. Given the BNNecks module , is the logit vector.
@@
where denotes either weights or activations of a specific layer, returns with values below set to and values above set to ( and are scopes set based on quantization type and bit-width), is a learnable parameter that adjusts the quantization step size and is the round function. For example, given a quantization bit-width , and is set to and respectively for unsigned input and and for signed input. Since the round function is non-differentiable, the derivative of with respect to during the backward pass [9; 6; 45] can be represented as:
@@ 3.2 Training with Soft Quantizer
STE serves as a classical method for estimating gradients, proven to be simple yet effective in many tasks. For some learning-based quantization methods, using STE to estimate quantization errors has yielded promising results. This is because, during quantization-aware training, algorithms tend to minimize task loss rather than quantization error [9; 25]. However, in the context of gait recognition tasks, particularly in appearance-based approaches with binarized input, even minor numerical perturbations can lead to vastly different features. Therefore, it should be helpful to use a differentiable function to approximate the step function to enhance the performance of gait recognition tasks. For the round function, the maximum quantization error occurs at the midpoint between two adjacent approximate values. Additionally, we hope that the derivative of this function varies periodically with changes in input values. A function that meets the above requirements can be defined as follows:
@@
Assuming the random variable follows a uniform distribution, we can compute the first term as (see Appendix A.1):
which is a function involving that is consistently less than as Fig. 3(b) shows. As increases, this term tends to approach , indicating that the optimization process will become increasingly challenging to continue. But we aim for as much descent as possible at each iteration, so STE is better for the first term. As for the second term, we can compute that:
which is consistently greater than as Fig shows. Combining Eq. 8, Eq. 11, and Eq. 12, we can see that while the soft quantizer simulates quantization error, STE is superior in terms of convergence as Fig. 3(c) shows. Therefore, during the initial training phase, STE is utilized to minimize task loss, while during the second fine-tuning phase, the soft quantizer is employed to balance quantization error and task loss.
3.3 Inter-class Distance-guided Distillation
In the preceding sections, we mitigated the impact of quantization error on prediction results by introducing the soft quantizer. However, on some complex datasets[33; 43], transitioning from 32-bit to 8-bit quantization still inevitably incurs some loss due to inherent information loss in the process of data discretization. Moreover, we observed that for the embeddings outputted by the 4-bit network, the relative distance between samples of different classes has changed. This inspires us to explore whether we can leverage information from high-bit networks to guide the optimization process of low-bit networks.
Given an input , the output embedding vector is . For with high bit and low bit , the logits vectors are denoted as and respectively. Traditional Knowledge Distillation (KD) approaches measure the distribution difference between two vectors. The probability vectors are calculated as:

### v2

@@ Abstract
Existing deep learning methods have made significant progress in gait recognition. Quantization can facilitate the application of gait models as a model-agnostic general compression technique. Typically, appearance-based models binarize inputs into silhouette sequences. However, mainstream quantization methods prioritize minimizing task loss over quantization error, which is detrimental to gait recognition with binarized inputs. To address this, we propose a differentiable soft quantizer, which better simulates the gradient of the round function during backpropagation. This enables the network to learn from subtle input perturbations. However, our theoretical analysis and empirical studies reveal that directly applying the soft quantizer can hinder network convergence. We addressed this issue by adopting a two-stage training strategy, introducing a soft quantizer during the fine-tuning phase. However, in the first stage of training, we observed a significant change in the output distribution of different samples in the feature space compared to the full-precision network. It is this change that led to a loss in performance. Based on this, we propose an Inter-class Distance-guided Calibration (IDC) strategy to preserve the relative distance between the embeddings of samples with different labels. Extensive experiments validate the effectiveness of our approach, demonstrating state-of-the-art accuracy across various settings and datasets. Code is available at link.
@@ 1 Introduction
In recent years, gait recognition has emerged as a prominent field due to its potential in long-range pedestrian retrieval based on walking patterns. Unlike other biometric technologies, gait, as a form of locomotion, is difficult to disguise [Lam2007HumanGR, Connor2018BiometricRB]. Additionally, gait recognition is robust against common covariates such as attire, carrying items, and standing conditions [Guan2015OnRT, Choudhury2016ClothingAC]. Therefore, gait recognition methods hold significant promise for capturing biometric features remotely in uncontrolled environments, thus offering broad application prospects.
@@
With the development of deep learning, gait recognition methods based on deep features [Wu2017ACS, Makihara2017JointIA, Fan2020GaitPartTP, Hou2020GaitLN, Lin2020GaitRV, Lin2020GaitRW, guo2025gait, wang2025gaitadapt] have achieved notable results on various datasets [Yu2006AFF, Takemura2018MultiviewLP, Zheng2022GaitRI, Shen2022LidarGaitB3, Zhu2021GaitRI, Fan2022OpenGaitRG, Fan2023ExploringDM]. Currently, state-of-the-art gait recognition technologies can achieve recognition distances exceeding hundreds of meters even with 4K high-definition cameras [Shen2022LidarGaitB3]. However, existing methods have scarcely considered the increasing computational resources required by the advanced edge devices. In real-world gait recognition scenarios, especially when real-time processing of video streams is required, the inference overhead of gait models cannot be overlooked as Fig. 1 shows. While methods based on CNNs [Fan2022OpenGaitRG, Zheng2022GaitRI] or Transformers [Fan2023ExploringDM] have demonstrated excellent performance, edge devices in deployment environments cannot meet the demands of these resource-intensive algorithms. This necessitates reducing the memory and computational burden of gait recognition methods while preserving performance, enabling deployment on resource-constrained devices. Current efforts to compress and accelerate neural networks primarily focus on quantization [Zhou2016DoReFaNetTL, Choi2018PACTPC, Esser2019LearnedSS, Lee2021NetworkQW] and pruning [Lin2020HRankFP, Zhou2021LearningNM, He2018FilterPV]. Compared to pruning, quantization is more deployment-friendly and offers significant acceleration, particularly Quantization Aware Training [Jacob2017QuantizationAT].
@@
Existing quantization methods [Zhang2018LQNetsLQ, Baskin2018NICENI, Esser2019LearnedSS] have achieved decent results on some tasks, but applying them to gait recognition tasks has failed to deliver comparable performance. These methods use a Straight-Through Estimator (STE) [Bengio2013EstimatingOP] to approximate the gradient of non-differentiable operations during the backward pass. While STE can smoothly transfer gradients, it is not sensitive to decimal-level changes. This may have no impact on other tasks but can be fatal for changes in gait posture. For example, in RGB input, the variation from 0 to 1 in a small area is not significant. However, for binarized gait silhouettes, this may indicate a change in posture. In this paper, we propose using parameterized soft quantization operators to prevent the network from accumulating quantization errors. Initially, we attempted to directly utilize a soft quantizer in the training process as a replacement for STE. However, through theoretical analysis and empirical validation, we have demonstrated that although this approach can simulate the quantization error introduced by non-differentiable operations, it significantly impacts the network’s convergence. To harness the strengths of both STE and the soft quantizer while mitigating their respective limitations, we propose a two-stage training strategy. In the first stage, we employ STE to facilitate the network’s convergence towards the vicinity of the optimal solution. Subsequently, we utilize the soft quantizer in the fine-tuning stage, enabling the simulation of decimal-level errors, which improves the network performance.
@@
After quantization, the 8-bit models exhibit outstanding performance, surpassing even the full-precision model. However, when quantized to lower bit widths, the model’s accuracy experiences significant degradation. Through analysis of the sample distribution in feature space, we observed significant changes in the relative positions of samples between the 4-bit quantized model and the full-precision model. This change could be a significant contributing factor to the observed decrease in accuracy. This insight prompts us to consider whether a calibration strategy could be employed to align the sample distribution of the low-bit models to high-bit models. A possible alignment solution is to distill the distribution from high-bit models. However, through an examination of the properties of naive distillation [Esser2019LearnedSS, Kim2019QKDQK, Yao2022ZeroQuantEA], we found that such methods result in the low-bit model learning the numerical discrepancies between the outputs of the full-precision model and itself, which are discrepancies inherently introduced by quantization. Therefore, we propose an Inter-class Distance-guided Calibration strategy to focus more on the differences in sample distribution between classes rather than the numerical differences within classes.
@@
We introduce a quantization-based gait recognition network (QGait), which performs lossless compression and acceleration on existing methods as Fig. 2 shows.
We propose a two-stage quantization training strategy together with the IDC method, which simultaneously optimizes quantization error and task loss.
@@ 2.1 Gait Recognition
Gait recognition methods can generally be classified into two categories: model-based methods [Li2020EndtoEndMG, Liao2017PoseBasedTN] and appearance-based methods [Wang2023DyGaitED, Shen2022GaitRW]. Model-based approaches tend to utilize estimated human body models as input, such as 2D/3D poses and SMPL [Loper2023SMPLAS]. For instance, GaitGraph [Teepe2021GaitgraphGC] employs graph convolutional networks for learning gait representations based on 2D skeletons, SMPLGait [Zheng2022GaitRI] leverages the 3D geometric information of SMPL models to enhance the learning of gait appearance features, and GaitTR [Zhang2022SpatialTN] combines transformer and convolutional layers to represent spatial and temporal information respectively. However, although model-based methods theoretically exhibit robustness to factors like carrying items and clothing, they often perform poorly due to the error accumulation in the pre-processing stage while entailing high computational costs, thereby potentially lacking practicality in certain real-world scenarios. On the other hand, appearance-based methods directly learn shape features from input videos, offering simplicity and ease of use while preserving privacy. With the rise of deep learning, most current appearance-based works focus on spatial feature extraction and gait temporal modeling. Specifically, GaitSet [Chao2018GaitSetRG] treats gait sequences as a set for the first time, employing the maximum function to compress sequences of frame-level spatial features. GaitPart [Fan2020GaitPartTP] meticulously explores minor differences in various parts of the silhouette to model periodic gait features. GaitGL [Lin2020GaitRV] addresses the limitations of spatially global gait representations in neglecting details, and locally based descriptors in failing to capture relationships between adjacent parts, by developing global and local convolutional layers. Building upon these works, GaitBase [Fan2022OpenGaitRG] significantly simplifies gait modeling and achieves excellent results with its simple yet effective design. Considering that existing methods often lack a practical perspective, it becomes necessary to design a lightweight network to aid gait recognition models in real-world applications. To the best of our knowledge, the proposed QGait is the first quantization-based compressed gait recognition network.
@@ 2.2 Network Quantization
Quantization [Chen2024AdaptiveQW, Cai2023BinarizedSC, Rastegari2016XNORNetIC] is a widely adopted technique in computer vision for compressing and accelerating neural networks. The advantage of quantization is that it can be easily combined with pruning, distillation, and other methods to further compress the network. Depending on the mapping strategy, quantization can be categorized into uniform and non-uniform [Liu2021NonuniformtoUniformQT] quantization. Non-uniform quantization necessitates specific hardware support, thus current research predominantly focuses on uniform quantization exploration. Uniform quantization methods encompass Post-Training Quantization (PTQ) [Shang2022PostTrainingQO] and Quantization Aware Training (QAT). PTQ has gained popularity for its ability to quantize models without requiring retraining. However, it suffers from accuracy degradation due to its heavy reliance on input data for setting quantization parameters, leading to poor alignment between parameters and weights. QAT addresses this issue by training with inserted fake quantization nodes [Jacob2017QuantizationAT], allowing network weights to adapt to quantization errors and enhancing the accuracy of quantized models. Early works like Dorefa [Zhou2016DoReFaNetTL] achieve network acceleration by quantizing activation values, weights, and gradients simultaneously. Further advancements such as PACT [Choi2018PACTPC], utilize learnable parameters to clip activation values to gain more reasonable quantization intervals. LSQ [Esser2019LearnedSS], one of the most widely used quantization methods, introduces a novel approach to estimate and scale the task loss gradient on the quantizer step size for each weight and activation layer, enabling the quantization step size to be learned with network optimization and achieving commendable results across various tasks. Given the satisfactory accuracy and notable acceleration achieved by quantized networks, quantization-based model compression has been extensively studied in various domains [Tian2023CABMCB, Yao2022ZeroQuantEA, Lin2023AWQAW], demonstrating the efficacy of quantization schemes. However, the impact of quantization in gait recognition remains underexplored. In this paper, we introduce QGait, which, built upon optimized quantization training strategies and tailored to the characteristics of gait recognition tasks, unveils the effectiveness of quantization in compressing gait recognition networks for the first time, offering novel perspectives and insights into gait recognition research.
@@ 3.1 Preliminaries
where denotes the embeddings, is the separate fully connected layer, is the temporal pooling and horizontal pooling layer, is the backbone network, and denotes the connection among network parts. Given the BNNecks module , is the logits vector.
@@
where denotes either weights or activations of a specific layer, returns with values below set to and values above set to ( and are scopes set based on quantization type and bit-width), is a learnable parameter that adjusts the quantization step size and is the round function. For example, given a quantization bit-width , and is set to and respectively for unsigned input and and for signed input. Since the round function is non-differentiable, the derivative of with respect to during the backward pass [Esser2019LearnedSS, Choi2018PACTPC, Zhou2016DoReFaNetTL] can be represented as:
@@ 3.2 Training with Soft Quantizer
STE serves as a classical method for estimating gradients, proven to be simple yet effective in many tasks. For some learning-based quantization methods, using STE to estimate quantization errors has yielded promising results. This is because, during quantization-aware training, algorithms tend to minimize task loss rather than quantization error [Esser2019LearnedSS, Liu2021NonuniformtoUniformQT]. However, in the context of gait recognition tasks, particularly in appearance-based approaches with binarized input, even minor numerical perturbations can lead to vastly different features. Therefore, it should be helpful to use a differentiable function to approximate the step function to enhance the performance of gait recognition tasks. For the round function, the maximum quantization error occurs at the midpoint between two adjacent approximate values. Additionally, we hope that the derivative of this function varies periodically with changes in input values. A function that meets the above requirements can be defined as follows:
@@
Assuming the random variable follows a uniform distribution, we can compute the first term as:
which is a function involving that is consistently less than as Fig. 4(b) shows. As increases, this term tends to approach , indicating that the optimization process will become increasingly challenging to continue. However, we aim for as much descent as possible at each iteration, so STE is better for the first term. As for the second term, we can compute that:
which is consistently greater than . Combining Eq. 8, Eq. 11, and Eq. 12, we can see that while the soft quantizer simulates quantization error, STE is superior in terms of convergence as Fig. 4(c) shows. Therefore, during the initial training phase, STE is utilized to minimize task loss, while during the second fine-tuning phase, the soft quantizer is employed to balance quantization error and task loss.
3.3 Inter-class Distance-guided Calibration
| GaitSet [Chao2018GaitSetRG] | 32 | 10.68 | 36.75 | 58.66 | 64.20 | 30.21 | 17.55 |
| GaitPart [Fan2020GaitPartTP] | 32 | 10.67 | 28.44 | 47.58 | 53.70 | 22.01 | 12.32 |
| GaitGL [Lin2020GaitRV] | 32 | 41.64 | 29.88 | 48.69 | 54.42 | 22.30 | 13.50 |
| DyGait [Wang2023DyGaitED] | 32 | 652.22 | 66.30 | 80.80 | 86.10 | 56.40 | 37.30 |
| GaitBase [Fan2022OpenGaitRG] | 32 | 118.30 | 64.40 | 81.50 | 85.80 | 55.28 | 36.73 |
| Dorefa-4 [Zhou2016DoReFaNetTL] | w4/a4 | - | 60.00 | 78.30 | 83.20 | 50.32 | 32.66 |
| PACT-4 [Choi2018PACTPC] | w4/a4 | - | 59.80 | 78.00 | 82.80 | 49.56 | 31.53 |
| LSQ-4 [Esser2019LearnedSS] | w4/a4 | - | 63.80 | 79.60 | 86.30 | 54.22 | 35.71 |
| QGait-4 | w4/a4 | 1.85 | 64.60 | 81.20 | 86.20 | 54.29 | 35.76 |
| Dorefa-8 [Zhou2016DoReFaNetTL] | w8/a8 | - | 64.30 | 80.80 | 86.10 | 54.34 | 35.87 |
| PACT-8 [Choi2018PACTPC] | w8/a8 | - | 63.30 | 80.50 | 85.20 | 53.41 | 34.64 |
| LSQ-8 [Esser2019LearnedSS] | w8/a8 | - | 64.80 | 81.40 | 85.90 | 55.15 | 36.76 |
| QGait-8 | w8/a8 | 7.39 | 66.50 | 81.30 | 86.20 | 55.82 | 37.05 |
In the preceding sections, we mitigated the impact of quantization error on prediction results by introducing the soft quantizer. However, on some complex datasets [Takemura2018MultiviewLP, Zheng2022GaitRI], transitioning from 32-bit to 8-bit quantization still inevitably incurs some loss due to inherent information loss in the process of data discretization. Moreover, we observed that for the embeddings outputted by the 4-bit network, the relative distance between samples of different classes has changed. This inspires us to explore whether we can leverage information from high-bit networks to guide the optimization process of low-bit networks.
Given an input , the output embedding vector is . For with high bit and low bit , the logits vectors are denoted as and respectively. Naive Knowledge Distillation (KD) approaches measure the distribution difference between two vectors. The probability vectors are calculated as: